OpenStudy (anonymous):

Find the slope of the tangent line to the curve 3 x^2 - 2 xy + 1 y^3 = 72 at the point ( -3,3 ).

OpenStudy (anonymous):

what is the "1" doing there?

OpenStudy (anonymous):

its in the problem

OpenStudy (anonymous):

thats how they wrote it...

OpenStudy (anonymous):

with a "1" as in \(1\times y^3\) very odd indeed

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

start with \[6x-2y-xy'+3y^2y'=0\] replace x by -1 and y by 3 and solve for \(y'\) or else solve for \(y'\) using algebra first and then make the replacement. no matter

OpenStudy (anonymous):

ok lets see

OpenStudy (anonymous):

damn typo \[6x-2y-2xy'+3y^2y'=0\]

OpenStudy (anonymous):

the only calculus part of the problem is finding the derivative. do you get that part?

OpenStudy (anonymous):

i dont understand how u get the implicit differentiation....

OpenStudy (anonymous):

ok think of it this way: you are pretending y is a function of x, so think of y as \(y=f(x)\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so for example, the derivative of \(y^3\) is like taking the derivative of \(f^3(x)\) by the chain rule you get \(3f^2(x)f'(x)\) but you write \(3y^2y'\) which is more brief

OpenStudy (anonymous):

ok i see

OpenStudy (anonymous):

and for \(xy\) you treat it like \(xf(x)\) so you need the product rule, just as you would it if was \(x\sin(x)\)

OpenStudy (anonymous):

so what if u have 3xy^2?

OpenStudy (anonymous):

then you need both the product and the chain rule. suppose you had \(3x\sin^2(x)\)

OpenStudy (anonymous):

you would have to use the product rule, because it is a product, and you have to use the chain rule to find the derivative of \(\sin^2(x)\)

OpenStudy (anonymous):

so if you have \(3xy^2\) you think of it as \(3xf^2(x)\) and the derivative would be \[3(f^2(x)+x\times 2f(x)f'(x)\] by both the product and the chain rule but you just write \[3(y^2+2xy')\]

OpenStudy (anonymous):

ok it makes sense.....thanks...

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

actually there is a another more prosaic way of looking at this. got to the video on implicit diff here http://justmathtutoring.com/ to get a different take on it if you need more

OpenStudy (anonymous):

k thanks so much